# The dimension and Bose distance of some BCH codes of length $\frac{q^{m}-1}{\lambda}$

### v1

@@ IV Auxillary Lemmas
If $\frac{q^{h+1}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$, then we have $q^{h+1}\leq\lambda(\delta-1)<q^{\lfloor(2m-1)/3\rfloor+1}$. This implies that $1\leq k_{\delta}\leq\lfloor({2m-1})/{3}\rfloor-h.$ By applying Theorem 1, we can conclude from Assertions 3 and 4 that
@@
By now, we have already demonstrated that both (34) and (35) hold for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$ Next, we show that equation (18) holds. Notice that $\lambda(\delta-1)<q^{m-\lfloor{m}/{3}\rfloor}$. We can apply Theorem 2 and Lemma 1 to obtain
$|C_{n}(a)|=|C_{\lambda n}(\lambda a)|=m\hbox{ or }\frac{m}{2}\quad\hbox{for any integer }a\in[1,\delta-1].$ (42)
@@
As examples of Theorem 3, we present the dimension of $\mathcal{C}_{(q,n,\delta)}$ for $2\leq\delta\leq\frac{q^{\lceil\frac{m}{2}\rceil+1}-1}{\lambda}+1$ by providing the following two Corollaries. Notably, the formulas given in Corollary 2 provide the dimension of the same narrow-sense BCH codes as those studied in [40, Theorem 3], but they are presented in a simpler form.
@@
If $2\leq\delta\leq\frac{q^{h}-1}{\lambda}+1$, then we have $\widetilde{f}(\delta)=0$ and $g(\delta)=0$. If $\delta>\frac{q^{h}-1}{\lambda}+1$, then we distinguish the following two cases.
@@
If $\delta\leq\frac{q^{h+1}-1}{\lambda}+1,$ we can directly have $f(\delta)=0$. If $\delta>\frac{q^{h+1}-1}{\lambda}+1,$ then we have $k_{\delta}=1$. We distinguish following cases.
Case 1. Suppose that $\delta_{h}>0$. Then we have $s_{\delta}=0$. This leads to
@@
Recall equation (13). It follows that
@@
Since each integer $t\in\left[\delta_{h+1}q+1,\delta_{h+1}q+\delta_{h}-1\right]$ can be uniquely expressed as $t=\delta_{h+1}q+i$ with $i\in\left[1,\delta_{h}-1\right]$, we have
@@
VI The Bose distance of $\mathcal{C}_{(q,n,\delta)}$
In this section, we investigate the Bose distance of BCH codes $\mathcal{C}_{(q,n,\delta)}$ for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$. Note that if $q\mid\delta$, then $\delta$ is not a coset leader modulo $n$. Therefore, we have $d_{B}(\mathcal{C}_{(q,n,\delta)})=d_{B}(\mathcal{C}_{(q,n,\delta+1)})$ for $q\mid\delta$. Given this property, it is sufficient to focus on BCH codes $\mathcal{C}_{(q,n,\delta)}$ with $q\nmid\delta$. Their Bose distances are established in the following three theorems.
@@ VI The Bose distance of $\mathcal{C}_{(q,n,\delta)}$
Let $m$ and $\delta$ be two positive integers with $2\leq\delta\leq\frac{q^{m-h}-1}{\lambda}$ and $q\nmid\delta$. Then $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta.$
@@
Note that $\lambda\mid q-1$ implies that $\mathrm{gcd}(q,\lambda)=1$. Since $q\nmid\delta$, it follows that $q\nmid\lambda\delta$. Additionally, it is clear that $2\leq\lambda\delta<q^{m-h}$. By [35, Theorem 2.3], it follows that $\lambda\delta$ is a coset leader modulo $\lambda n$. Then applying Lemma 1, we conclude that $\delta$ is a coset leader modulo $n$. Consequently, $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta.$
@@
Let $m\geq 5$ be an odd integer, and $\delta$ be an integer with $\frac{q^{h+1}-1}{\lambda}+1\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$ and $q\nmid\delta$. Let $j_{\delta}=\lfloor\log_{q}(\lambda\delta)\rfloor-h$ and $\sum\limits_{\ell=0}^{h+j_{\delta}}\delta_{\ell}q^{\ell}$ be the $q$-adic expansion of $\lambda\delta$. Let $r_{\delta}$ be the smallest integer in $j_{\delta}+1,\,j_{\delta}]$ such that $\delta_{h+r_{\delta}}>0$ and define $\hat{\delta}=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}.$ Then
@@
Case 1. Suppose that $\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}$. By Remarks 4 and 5, the inequality implies that $\lambda\delta\not\in\mathcal{A}_{j_{\delta}}$ for every integer $i\in[-j_{\delta}+1,j_{\delta}]$. Therefore, by Theorem 1, we conclude $\lambda\delta\not\in\mathcal{S}$. Note that the assumptions $q\nmid\delta$ and $\lambda\mid q-1$ imply $q\nmid\lambda\delta$. It follows that $\lambda\delta$ is a coset leader modulo $\lambda n$. Applying Lemma 1, we further obtain that $\delta$ is a coset leader modulo $n$. Consequently, we have $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta$.
Case 2. Suppose that $\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}$. With an argument similar to that in the proof of [39, Theorem 6], we can show that each integer in $\lambda\delta,\hat{\delta}$] is not a coset leader modulo $\lambda n$. Applying Lemma 1, it follows that each integer in $\delta,\lfloor\frac{\hat{\delta}}{\lambda}\rfloor]$ is not a coset leader modulo $n$. Notice that when $m$ is odd,
@@
Let $m\geq 4$ be an even integer, and let $\delta$ be an integer with $\frac{q^{h}-1}{\lambda}+1\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$ and $q\nmid\delta$. Let $j_{\delta}=\lfloor\log_{q}(\lambda\delta)\rfloor-h$ and $\sum\limits_{\ell=0}^{h+j_{\delta}}\delta_{\ell}q^{\ell}$ be the $q$-adic expansion of $\lambda\delta$. Let $r_{\delta}$ be the smallest integer in $j_{\delta},j_{\delta}]$ such that $\delta_{h+r_{\delta}}>0$ and define $\hat{\delta}=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})}.$
@@
We can use an argument analous to that in the proof of Theorem 5 to conclude that 52) hlods if $r_{\delta}\neq 0$. Thus, we now only demonstrate equation (53) by considering the following cases.
Case 1. Suppose that $\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}\geq\sum\limits_{\ell=h}^{h+j_{\delta}}\delta_{\ell}q^{\ell-h}$. If the inequality is strict, by Theorem 1 together and Remarks 45, we have $\lambda\delta\notin\mathcal{S}\cup\mathcal{H}$; If equality holds, by Lemma 2, we $\lambda\delta\in\mathcal{H}$. Note that the assumptions $q\nmid\delta$ and $\lambda\mid q-1$ imply $q\nmid\lambda\delta$. Therefore, in either case, $\lambda\delta$ is a coset leader modulo $\lambda n$. By Lemma 1, it follows that $\delta$ is a coset leader of modulo $n$. Therefore, $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta$.
@@
An $[n,k,d]$ linear code over $\mathbb{F}_{q}$ is called optimal if there does not exist an $n,k^{\prime},d]$ linear code with $k^{\prime}>k$, or an $n,k,d^{\prime}]$ linear code with $d^{\prime}>d$ over $\mathbb{F}_{q}$. Using the formulas provided in Theorems 3 and 6, we identify several narrow-sense BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ that are either optimal or have parameters matching those of the best known linear codes in the database. Most of these codes occur in the primitive case $\lambda=1$. See Tables II – III in our previous work [39] for examples of these primitive BCH codes. In addition, we list further examples of narrow-sense BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ with $\lambda\neq 1$ in Table I. The parameters of these codes are nearly optimal or best known. All parameters were verified by Magma, and for all BCH codes in the table, the Bose distance $d_{B}$ coincides with the minimum distance $d$.
Table I: Examples of BCH code $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ with $\lambda\neq 1$
@@
VII BCH codes $\mathcal{C}_{(q,n,\delta)}$with $\lambda\delta=aq^{h+k}+b$.
As an illustration of our main results, this section presents the dimension and Bose distance of the BCH code $\mathcal{C}_{(q,n,\delta)}$ with $\lambda\delta=aq^{h+k}+b$, where $m-2h\leq k\leq\lfloor(2m-1)/3\rfloor-h$, $1\leq a\leq q-1$, $\lambda\leq b\leq q^{m-h-k}$ and $q\nmid b$.
@@
Let $m\geq 5$ be an odd integer, and let $k$ be an integer with $1\leq k\leq\lfloor(2m-1)/3\rfloor-h$. If $\lambda\delta=aq^{h+k}+b$ for some integers $a\in[1,q-1]$ and $b\in[\lambda,q^{h-k+1}]$ such that $q\nmid b$, then
$d_{B}(\mathcal{C}_{(q,n,\delta)})=\begin{cases}\delta&\hbox{if }b>aq^{2k-1},\\ \lfloor\frac{aq^{h+k}+aq^{2k-1}}{\lambda}\rfloor+1&\hbox{if }b\leq aq^{2k-1}\end{cases}$ (54)
@@
Note that $q\nmid b$ implies that $q\nmid\lambda\delta$. Since $\lambda$ is a factor q1q-1, it follows that $q\nmid\delta$. Then by applying Theorem 5, we can directly derive equation (54).
@@
with $\sum\limits_{\ell=0}^{h-k}\delta_{\ell}q^{\ell}=b-\lambda$. By definition, it follows that
@@
Let $m\geq 4$ be an even integer, and let $k$ be an integer with $0\leq k\leq\lfloor(2m-1)/3\rfloor-h$. Suppose that $\lambda\delta=aq^{h+k}+b$ for some integers $a\in[1,q-1]$ and $b\in[\lambda,q^{h-k}]$ such that $q\nmid b$.
@@
If $k=0$, then we have $\lambda(\delta-1)=aq^{h}+b-\lambda$ with $0\leq b-\lambda<q^{h-1}$. It follows that
@@
with $\sum\limits_{\ell=0}^{h-k}\delta_{\ell}q^{\ell}=b-\lambda$. By definition, it follows that
@@
Then applying Lemmas 8 and 10, we obtain
@@
VIII Some non-narrow-sense BCH codes $\mathcal{C}_{(q,n,\delta,b)}$
In this section, we determine the Bose distance and dimension of certain non-narrow-sense BCH codes $\mathcal{C}_{(q,n,\delta,b)}$.

### v2

@@ V The dimension of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$
If $\frac{q^{h+1}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$, then we have $q^{h+1}\leq\lambda(\delta-1)<q^{\lfloor(2m-1)/3\rfloor+1}$. In this case, we have $1\leq k_{\delta}\leq\lfloor({2m-1})/{3}\rfloor-h.$ By applying Lemma 3, we can conclude from Assertions 3 and 4 that
@@
By now, we have already demonstrated that both (36) and (37) hold for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$ Next, we show that equation (20) holds. Notice that $\lambda(\delta-1)<q^{m-\lfloor{m}/{3}\rfloor}$. We can apply Lemmas 1 and 4 to obtain
$|C_{n}(a)|=|C_{\lambda n}(\lambda a)|=m\hbox{ or }\frac{m}{2}\quad\hbox{for all integers }a\in[1,\delta-1].$ (44)
@@
Zhu et al. determined the dimension of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$ for designed distance $2\leq\delta\leq\frac{q^{\lceil(m+1)/2\rceil}-1}{\lambda}+1$ in [41, Theorem 1] for odd $m$ and in [41, Theorem 3] for even $m$, respectively. As an illustration of Theorem 1 in this paper, we present the dimension of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$ for designed distances $2\leq\delta\leq\frac{q^{\lceil m/2\rceil+1}-1}{\lambda}+1$ in the following two corollaries. Notably, for even $m$, although the range of designed distances considered in Corollary 2 coincides with that in [41, Theorem 3], our derivation leads to a simpler expression, involving only four cases compared to eleven cases in [41, Theorem 3].
@@
If $2\leq\delta\leq\frac{q^{h}-1}{\lambda}+1$, then we have $\widetilde{f}(\delta)=0$ and $g(\delta)=0$.
If $\delta>\frac{q^{h}-1}{\lambda}+1$, then we distinguish the following two cases.
@@
If $\delta\leq\frac{q^{h+1}-1}{\lambda}+1,$ we directly have $f(\delta)=0$.
If $\delta>\frac{q^{h+1}-1}{\lambda}+1,$ then we have $k_{\delta}=1$. We distinguish following cases.
Case 1. Suppose that $\delta_{h}>0$. In this case, we have $s_{\delta}=0$. This leads to
@@
By substituting these values into equation (15), we obtain
@@
Since each integer $t\in\left[\delta_{h+1}q+1,\delta_{h+1}q+\delta_{h}-1\right]$ can be uniquely expressed as $t=\delta_{h+1}q+i$ with an integer $i\in\left[1,\delta_{h}-1\right]$, we have
@@
VI The Bose distance of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$
In this section, we investigate the Bose distance of narrow-sense BCH codes of length $n=(q^{m}-1)/\lambda$ for designed distances $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$. Note that if $q\mid\delta$, then $\delta$ is not a coset leader modulo $n$. Therefore, we have $d_{B}(\mathcal{C}_{(q,n,\delta)})=d_{B}(\mathcal{C}_{(q,n,\delta+1)})$ for $q\mid\delta$. Given this property, it is sufficient to focus on BCH codes $\mathcal{C}_{(q,n,\delta)}$ with $q\nmid\delta$. Their Bose distances are established in the following three theorems.
@@ VI The Bose distance of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$
Let $m$ be a positive integer, and let $\lambda$ be a positive divisor of q1q-1. Set $n=(q^{m}-1)/\lambda$ and $h=\lfloor\frac{m}{2}\rfloor$. Let $\delta$ be an integer with $2\leq\delta\leq\frac{q^{m-h}-1}{\lambda}$ and $q\nmid\delta$. Then $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta.$
@@
Note that $\lambda\mid q-1$ implies that $\mathrm{gcd}(q,\lambda)=1$. Since $q\nmid\delta$, it follows that $q\nmid\lambda\delta$. Additionally, it is clear that $2\leq\lambda\delta<q^{m-h}$. By Lemma 2, it follows that $\lambda\delta$ is a coset leader modulo $\lambda n$. Then applying Lemma 1, we conclude that $\delta$ is a coset leader modulo $n$. Consequently, $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta.$
@@
Let $m\geq 5$ be an odd integer, and let $\lambda$ be a positive divisor of q1q-1. Set $n=(q^{m}-1)/\lambda$ and $h=\lfloor\frac{m}{2}\rfloor$. Let $\delta$ be an integer with $\frac{q^{h+1}-1}{\lambda}+1\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$ and $q\nmid\delta$. Define $j_{\delta}=\lfloor\log_{q}(\lambda\delta)\rfloor-h$, and write the $q$-adic expansion of $\lambda\delta$ as $\lambda\delta=\sum\limits_{\ell=0}^{h+j_{\delta}}\delta_{\ell}q^{\ell}$. Let $r_{\delta}$ be the smallest integer in $j_{\delta}+1,\,j_{\delta}]$ such that $\delta_{h+r_{\delta}}>0$ and let $\hat{\delta}=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}.$ Then
@@
Case 1. Suppose that $\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}$. By Remarks 4 and 5, the inequality implies that $\lambda\delta\not\in\mathcal{A}_{j_{\delta}}(i)$ for every integer $i\in[-j_{\delta}+1,j_{\delta}]$. Therefore, by Lemma 3, we conclude $\lambda\delta\not\in\mathcal{S}$. Note that the assumptions $q\nmid\delta$ and $\lambda\mid q-1$ imply $q\nmid\lambda\delta$. It follows that $\lambda\delta$ is a coset leader modulo $\lambda n$. Applying Lemma 1, we further obtain that $\delta$ is a coset leader modulo $n$. Consequently, we have $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta$.
Case 2. Suppose that $\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}$. With an argument similar to that in the proof of [40, Theorem 6], we can show that each integer in $\lambda\delta,\hat{\delta}$] is not a coset leader modulo $\lambda n$. Applying Lemma 1, it follows that each integer in $\left[\delta,\lfloor\frac{\hat{\delta}}{\lambda}\rfloor\right]$ is not a coset leader modulo $n$. Notice that when $m$ is odd,
@@
Let $m\geq 4$ be an even integer, and let $\lambda$ be a positive divisor of q1q-1. Set $n=(q^{m}-1)/\lambda$ and $h=\lfloor\frac{m}{2}\rfloor$. Let $\delta$ be an integer such that $\frac{q^{h}-1}{\lambda}+1\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$ and $q\nmid\delta$. Define $j_{\delta}=\lfloor\log_{q}(\lambda\delta)\rfloor-h$, and write the $q$-adic expansion of $\lambda\delta$ as $\lambda\delta=\sum\limits_{\ell=0}^{h+j_{\delta}}\delta_{\ell}q^{\ell}$. Let $r_{\delta}$ be the smallest integer in $j_{\delta},j_{\delta}]$ such that $\delta_{h+r_{\delta}}>0$ and let $\hat{\delta}=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})}.$
@@
We can use an argument analogous to that in the proof of Theorem 3 to conclude that 54) holds if $r_{\delta}\neq 0$. Thus, we now only demonstrate equation (55) by considering the following cases.
Case 1. Suppose that $\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}\geq\sum\limits_{\ell=h}^{h+j_{\delta}}\delta_{\ell}q^{\ell-h}$. If the inequality is strict, by Lemma 3 and Remarks 45, we have $\lambda\delta\notin\mathcal{S}\cup\mathcal{H}$; If equality holds, by Lemma 5, we $\lambda\delta\in\mathcal{H}$. Note that the assumptions $q\nmid\delta$ and $\lambda\mid q-1$ imply $q\nmid\lambda\delta$. Therefore, in either case, $\lambda\delta$ is a coset leader modulo $\lambda n$. By Lemma 1, it follows that $\delta$ is a coset leader of modulo $n$. Therefore, $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta$.
@@
An $[n,k,d]$ linear code over $\mathbb{F}_{q}$ is called optimal if there does not exist an $n,k^{\prime},d]$ linear code with $k^{\prime}>k$, or an $n,k,d^{\prime}]$ linear code with $d^{\prime}>d$ over $\mathbb{F}_{q}$. Using the formulas given in Theorems 14, we can readily determine the dimension and Bose distance of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$ for designed distances $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$, where $\lambda$ is a positive divisor of q1q-1. By comparing these codes with the Database, we identify several BCH codes that are either optimal or have parameters matching those of the best-known linear codes. But most of these codes arise in the primitive case $\lambda=1$. See Tables II–III in our previous work [40] for examples of such primitive BCH codes.
In addition, we list further examples of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$ with $\lambda\neq 1$ in Table II. The parameters of these codes are nearly optimal or best-known in the Database. All parameters were verified by Magma, and for all BCH codes listed in Table II, the Bose distance $d_{B}$ coincides with the minimum distance $d$. The Python programs implementing these formulas, together with additional examples, are available at https://zheng-run.github.io.
Table II: Examples of BCH code $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ with $\lambda\neq 1$
@@
VII BCH codes with designed distance $\delta$ satisfying $\lambda\delta=aq^{h+k}+b$.
As an illustration of Theorems 14 established in the previous two sections, we derive the following two corollaries in this section. These corollaries present the dimension and Bose distance of narrow-sense BCH code $\mathcal{C}_{(q,n,\delta)}$ for $\delta$ satisfying $\lambda\delta=aq^{h+k}+b$, where $k,a,b$ are integers such that $m-2h\leq k\leq\lfloor(2m-1)/3\rfloor-h$, $1\leq a\leq q-1$, $\lambda\leq b\leq q^{m-h-k}$ and $q\nmid b$.
@@
Let $m\geq 5$ be an odd integer. Let $\delta$ be an integer such that $\lambda\delta=aq^{h+k}+b$ for some integers $k$, $a$, $b$ with $1\leq k\leq\lfloor(2m-1)/3\rfloor-h$, $1\leq a\leq q-1$, $\lambda\leq b\leq q^{h-k+1}$ and $q\nmid b$. Then
$d_{B}(\mathcal{C}_{(q,n,\delta)})=\begin{cases}\delta&\hbox{if }b>aq^{2k-1},\\ \lfloor\frac{aq^{h+k}+aq^{2k-1}}{\lambda}\rfloor+1&\hbox{if }b\leq aq^{2k-1},\end{cases}$ (56)
@@
Note that $q\nmid b$ implies that $q\nmid\lambda\delta$. Since $\lambda$ is a divisor of q1q-1, it follows that $q\nmid\delta$. Then by applying Theorem 3, we can directly derive equation (56).
@@
with $\sum\limits_{\ell=0}^{h-k}\delta_{\ell}q^{\ell}=b-\lambda$. In this case, we have
@@
Remark 6.
In Corollary 4, we assume $q\nmid b$. Since $\lambda$ is a divisor of q1q-1, we have $\gcd(q,\lambda)=1$. From $\lambda\delta=aq^{h+k}+b$, it follows that $q\mid b$ if and only if $q\mid\lambda\delta$, which is further equivalent to $q\mid\delta$. In this case, we have $\mathcal{C}_{(q,n,\delta)}=\mathcal{C}_{(q,n,\delta+1)},$ while clearly $q\nmid(\delta+1)$. Hence it suffices to consider the case $q\nmid b$. Similarly, in Corollary 5 below we also impose the condition $q\nmid b$.
@@
Let $m\geq 4$ be an even integer. Let $\delta$ be an integer such that $\lambda\delta=aq^{h+k}+b$ for some integers $k,a,b$ with $0\leq k\leq\lfloor(2m-1)/3\rfloor-h$, $1\leq a\leq q-1$, $\lambda\leq b\leq q^{h-k}$ and $q\nmid b$.
@@
If $k=0$, then we have $\lambda(\delta-1)=aq^{h}+b-\lambda$ with $0\leq b-\lambda<q^{h-1}$. In this case, we have
@@
with $\sum\limits_{\ell=0}^{h-k}\delta_{\ell}q^{\ell}=b-\lambda$. In this case, we have
@@
Then applying Lemmas 11 and 13, we obtain
@@
VIII Some non-narrow-sense BCH codes of length $(q^{m}-1)/\lambda$
In the previous sections, we study the parameters of narrow-sense BCH codes of length $n=(q^{m}-1)/\lambda$. For the dimension of narrow-sense BCH codes $\mathcal{C}_{(q,n,\delta)}$, the equality in (4) allows us to compute $\dim(\mathcal{C}_{(q,n,\delta)})$ by determining the coset leaders modulo $n$ within the range $1,\delta-1]$ and the sizes of their corresponding $q$-cyclotomic cosets. However, this approach cannot be directly applied to non-narrow-sense BCH codes $\mathcal{C}_{(q,n,\delta,b)}$ with $b\geq 2$, since we do not have an analogue of (4) for these codes. In general, for $b\geq 2$, we do not have
$\dim(\mathcal{C}_{(q,n,\delta,b)})=n-\sum_{a\in\mathcal{L}^{n}(b,b+\delta-2)}|C_{n}(a)|,$
because the equality $\left|\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(a)\right|=\sum\limits_{a\in\mathcal{L}^{n}(b,b+\delta-2)}|C_{n}(a)|$ may not hold when $b\geq 2$. Consequently, determining the dimension of non-narrow-sense BCH codes $\mathcal{C}_{(q,n,\delta,b)}$ with $b\geq 2$ is in general more difficult, and there are very few results on the parameters of such codes in the literature.
Nevertheless, although the above equality does not hold in general, the results on the coset leaders and sizes of $q$-cyclotomic cosets given in Sections II and III enable us to determine the dimension and Bose distance of certain classes of non-narrow-sense BCH codes, as shown in the following result.